A car traveling at 95km/h overtakes a 1.10km Zambia Railways train traveling in the same direction on a track parallel to the road. The train’s speed is 75km/h. (a) How long does it take the car to pass the train? (b) How far will the car have traveled in this time?
1
Expert's answer
2021-03-29T09:00:25-0400

a)


"t=\frac{l}{v}\\t=\frac{1.1}{95-75}=0.055\ h=3.3\ min"

b)


"s=(95)(0.055)=5.225\ km=5225\ m"